# Motion Primitives Planning For Center-Articulated Vehicles

### v1

@@ Abstract
Autonomous navigation across unstructured terrains, including forests and construction areas, faces unique challenges due to intricate obstacles and the element of the unknown. Lacking pre-existing maps, these scenarios necessitate a motion planning approach that combines agility with efficiency. Critically, it must also incorporate the robot’s kinematic constraints to navigate more effectively through complex environments. This work introduces a novel planning method for center-articulated vehicles (CAV), leveraging motion primitives within a receding horizon planning framework using onboard sensing. The approach commences with the offline creation of motion primitives, generated through forward simulations that reflect the distinct kinematic model of center-articulated vehicles. These primitives undergo evaluation through a heuristic-based scoring function, facilitating the selection of the most suitable path for real-time navigation. To augment this planning process, we develop a pose-stabilizing controller, tailored to the kinematic specifications of center-articulated vehicles. During experiments, our method demonstrates a $67\%$ improvement in SPL (Success Rate weighted by Path Length) performance over existing strategies. Furthermore, its efficacy was validated through real-world experiments conducted with a tree harvester vehicle - SAHA.
@@ I Introduction
Traditional planning methods [4], including sampling-based [5], grid-search [6], and optimization-based strategies [7], face challenges in real-world applications, particularly within unknown environments. Those environments are characterized by frequent and unpredictable changes, making it infeasible to maintain an updated global map which sampling-based methods and grid-search methods primarily rely on. Besides, the special and complex kinematic models of the vehicles make it difficult to find solutions efficiently through optimization-based methods in real-time operations. Nevertheless, end-to-end learning methods, despite directly mapping sensory observations into trajectories or actions, struggle to incorporate hard kinematic constraints into the network, leading to potential issues in path feasibility [8].
Figure 1: Experiment of a center-articulated vehicle navigating through a construction site. A, B, and C represent three intermediate moments. The goal is set in the orange dot. The green curve shows the vehicle’s trajectory avoiding obstacles marked by red boxes. The bottom images illustrate how obstacles influence the planning choices.
@@
This paper addresses the challenge of autonomous motion planning for center-articulated vehicles by introducing an algorithm that efficiently generates motion primitives tailored to their specific kinematics, combined with a receding horizon strategy for adaptive, real-time planning and a pose-stabilizing controller to ensure target convergence within kinematic and control constraints. The main contributions of this paper are summarized as follows:
@@
    A pose-stabilizing controller designed to ensure the convergence to a given target under the specific kinematic constraints and control limits of center-articulated vehicles.
@@ II Related Work
The kinematics and dynamics of CAVs have been studied [29, 30] to improve their autonomy in challenging environments, yet current planning research primarily leverages traditional methods. Nayl and Thaker’s optimization-based Model Predictive Control (MPC) approach aligns well with kinematics but lacks online efficiency [31]. Leander Peter’s use of Dubins curves ensures path traceability but lacks agility in obstacle-rich environments [32]. For these planning methods, various path-following control approaches were employed: Delrobaei and Mehdi’s Lyapunov-based approach uses polar coordinates to track the target point but does not inspect the trajectory [33], Altafini proposes a Frenet frame-based path-tracking control [34], however, the model is complicated. Ridley and Cork’s method [35] shows the theoretical feasibility of autonomous regulation but its linearized model is constrained by the assumption of a small steer angle. A modified pure-pursuit algorithm has been proposed and successfully applied to tractors and rovers offering the advantages of simplicity and stability [36, 37], yet the steering and speed control are still coupled. While these studies address specific problems, a systematic planning and control method with efficiency and agility needed for real-time, obstacle-rich environments is yet to be seen on CAVs.
@@ III Methodology
This section is structured as follows: III-A derived the kinematic model for CAVs, establishing the relationship between velocity (both linear and angular) and vehicle state variables, which is essential for the offline primitives generation introduced in III-B. Then, the optimal path group is selected from the primitive sets through an online receding horizon planner described in III-C. Finally, the path-following controller proposed in III-D samples a target point from the optimal path group for precise trajectory tracking.
@@ III-A Modeling of The Center-Articulated Vehicles
The schematic diagram of a CAV steering at a fixed steering angle is shown in Fig. 2. The distances from the central joint to the centers of wheels $P_{1}$ and $P_{2}$ of the front and rear parts are $l_{1}$ and $l_{2}$ respectively. $\gamma$ is the center steering angle and the orientation angles of the front and rear parts are $\theta_{1}$ and $\theta_{2}$ respectively.
@@
We define the velocity of the center of the front part $P_{1}$ as $v_{f}$, and the velocity of the base $P_{0}$ as $v_{b}$. (1) and (2), we get: As proven in [30], the relationship between the front part’s heading, velocity, steer angle and steer velocity can be expressed as:
@@
$\dot{\theta}_{1}=-\frac{v_{f}\sin(\gamma)+l_{2}\dot{\gamma}}{l_{1}\cos(\gamma)+l_{2}}$ (3)
The relationship of $v_{b}$ and $v_{f}$ can also be easily derived as:
$v_{b}\cos(\alpha)=v_{f}$ (4)
$v_{b}\sin(\alpha)=-l_{1}\dot{\theta}_{1}$ (5)
Considering the convenience for control, we focus on $v_{f}$ instead of $v_{b}$ since it is directly linked to the motor of the front part which can be controlled. Moreover, $v_{b}$ is jointly affected by $v_{f}$ and another control state variable: $\dot{\gamma}$, which is regulated by the torque at the center joint. To simplify the model and cater to the control variables $v_{f}$ and $\dot{\gamma}$, we set the front center $P_{1}$ as the virtual control point instead of $P_{0}$. This setting will be demonstrated to be convenient for the pose-stabilizing control law proposed in Section D. For simplicity, we will use $v$ to represent $v_{f}$ in subsequent discussions. Thus, our model simplifies to:
$\begin{bmatrix}\dot{x}_{1}\\ \dot{y}_{1}\\ \dot{\theta}_{1}\\ \dot{\gamma}\end{bmatrix}=\begin{bmatrix}\cos(\theta_{1})&0\\ \sin(\theta_{1})&0\\ -\frac{\sin(\gamma)}{L}&-\frac{l_{2}}{L}\\ 0&1\end{bmatrix}\begin{bmatrix}v\\ \dot{\gamma}\end{bmatrix}$ (6)
@@ III-B Offline Primitives Generation
The generation of primitives refers to the generation methods of state lattice primitives and control sampling primitives. According to the kinematic model in Equation (6), the vehicle’s state space includes $x_{1}$, $y_{1}$, $\theta_{1}$, $\gamma$, while the control space includes $v$ and $\dot{\gamma}$. First, we discretize the state space to create different groups of primitives. For the egocentric local planning problem, vehicle-centered coordinate frames can be used, which simplifies the vehicle’s state space as $(x_{1},y_{1})=(0,0)$ and $\theta_{1}=0$. Thus, we only discretize $\gamma$ into 30 groups, serving as the vehicle’s state lattices.
We define control groups $\Sigma_{ij}$ inside each state lattice $\Sigma_{i}$ as trajectories that vary from different initial control inputs. Each state lattice contains 15 control groups with different initial control inputs designed based on the mechanical limitations of vehicles and excluding significant overlap ($i\in[1,30],j\in[1,15]$). Each control group contains trajectories that start with the same $v$ and $\dot{\gamma}$ and split twice at $t=(3/v_{j})$s and $t=(6/v_{j})$s. Trajectories in each control group are generated using forward kinematics simulation:
@@ III-C Online Receding-horizon Planner
Primitives set generated in III-B are further processed for collision detection. Drawing inspiration from Zhang’s method [9], the proposed two-step collision detection method combines offline precomputation and online path selection. Initially, potential collision points along trajectories in primitive sets are identified and mapped to grid cells, which are later filled with real-time terrain data. During online running, the planner dynamically eliminates paths based on real-time obstacle data matched to these grids, ensuring efficient and responsive navigation without excessive computational demands. As shown in Fig. 4.
@@
Figure 4: An illustration of the two-step collision detection: (a) shows an example of marking the corresponding collision grid for the Trajectory $1$. (b) shows an online collision detection where Trajectory $3$ is marked as occluded.
@@
$S_{ij}=\frac{\sum_{k=1}^{N_{ij}}s^{ij}_{k}}{N_{ij}}$ (10)
Where each trajectory score $s^{ij}_{k}$ is determined by:
$\begin{split}s^{ij}_{k}=&(s^{dir}+\alpha\cdot s^{dist})^{2}\cdot s^{vel}\cdot s^{state}\cdot s^{terrain}\cdot s^{p}\end{split}$ (11)
in which each score term is given by:
$s^{dist}=$ $f^{dist}(dx,dy,D_{max})$ (12)
$s^{dir}=$ $f^{dir}({d\theta}_{1},{d\theta}_{2})$ (13)
$s^{state}=$ $f^{state}(\gamma_{k},\gamma_{now})$ (14)
$s^{vel}=$ $f^{vel}(v_{k},v_{max})$ (15)
@@
Following previous methods [9], we take the coinciding part before the first split ($3m$ length, as discussed in III-B) of the trajectories in the control group with the highest cumulative score, as the optimal path for the path-following controller to track. To sum up, the online local planning algorithm is described in Algorithm 1.
@@ III-D Pose-Stabilizing Control Law
The selected optimal path is accepted as input for the path-following controller. Considering the small time intervals of our receding horizon planner, we pick a target point from the chosen path instead of the whole path to track. After tuning, we set the look-ahead distance as 1.5m, which balances the responsiveness and the smoothness of trajectories. Adopting Park’s method [38], this control law is modified to accommodate the specialized kinematics of CAVs. It starts with the error definition in the polar coordinate. As shown in Fig. 2, the distance to the goal is $r$. Let $\phi\in(-\pi,\pi]$ represent the orientation of $P_{1}$ relative to the goal. Let $\delta\in(-\pi,\pi]$ denote the vehicle’s heading orientation relative to the lines toward the goal.
@@ III-D2 Fast Subsystem and Closed-Loop Steering
$\kappa=-\frac{1}{r}\left[k_{\delta}(\delta-\arctan(-k_{\phi}\phi))+\left(1+\frac{k_{\phi}}{1+(k_{\phi}\phi)^{2}}\right)\sin(\delta)\right]$ (24)
@@
where, $v_{op}$ is the operational speed and $v_{max}$ is the maximum linear velocity. $\beta$ and $\lambda$ are design parameters, chosen such that $v\rightarrow v_{op}$ as $\kappa\rightarrow 0$, and $v\rightarrow v_{max}$ as $\kappa\rightarrow\kappa_{max}$.
It should be noticed that the control law defined by Equations (25) to (27) is fundamentally reliant on pre-computed primitives for theoretical stability since both $\gamma$ and $\dot{\gamma}$ are mechanically constrained. When $\gamma$ or $\dot{\gamma}$ reaches the limit, Equation (25) might not be able to satisfy, resulting in potentially slow or even failure in the convergence of slow subsystem. This issue is solved by restricting each point in tracking paths: all points taken from the specifically designed primitive sets, which are tailored to CAVs’ kinematic and mechanical constraints, ensure the system’s convergence without encountering the outlined constraint.
@@ IV Experiments
We evaluated the proposed method in simulation environments and compared its performance with the baseline before carrying out real-world experiments. The test platform, SAHA (Supervised Autonomous HArvester) [39], is an autonomous tree harvester with a center-articulated chassis. The vehicle is equipped with a LiDAR for real-time environmental mapping and an IMU for chassis orientation data. Key specifications of SAHA are listed in Table I. Fig. 6 depicts the actual SAHA vehicle.
@@
TABLE II: Comparison of Motion Planning Test Results
@@ IV-A Simulation Experiments
The simulation experiment of SAHA is performed in the Gazebo simulator. We first evaluate the proposed path-following controller independently before using it in the autonomous driving pipeline. The controller is tested against the modified pure-pursuit approach, with the reference paths derived from generated primitives that accommodate different vehicle states. Specifically, three states are tested, each assessed across 90 trajectories. Cross-track error (CTE) [40] is used to evaluate the tracking error. As shown in Table III, our method’s advantage grows with the tracking steering angle $\gamma$ increases, showing a 4.73% reduction of CTE at $\gamma=0^{\circ}$ and 27% at $\gamma=30^{\circ}$.
TABLE III: CTE Comparison of Two Controller
@@
We then compared the proposed motion planning pipeline against the baseline method. The baseline uses a planner adapted from [9] with cropped cubic splines as primitives. These primitives, designed for searching within specified steering angles as depicted in Fig. 8, enhance the effectiveness of the search process to meet the vehicle’s kinematic constraints. The pure-pursuit (p.p.) controller is used in the baseline to track the path. In addition, we compared ours with the method that uses our primitives in the local planner, but without using our proposed pose-stabilizing controller for path following. All approaches utilize a geometric-based terrain analysis module [41] for obstacle detection without known prior maps.
@@
To evaluate robustness across various scenarios, we designed four environments with varying obstacle numbers $N_{t}$ (the number of tree obstacles) and $N_{w}$ (the number of wall obstacles). For each environment, we generated 2 random maps, each spanning $40m$ by $40m$ and featuring cylinder and box obstacles to represent trees and walls. 25 target points were randomly sampled on each map, totaling 200 points across all maps. As shown in Fig. 7, the green circles are tree obstacles, gray lines are wall obstacles and the red dots are goals.
The RRT* algorithm, based on Dubins Curve, constrained by the vehicle’s steer angle and minimum turning radius, was used to find approximated shortest paths to these points with given prior maps. During our experiments, we assume after 500,000 iterations, the solution could approximate the optimal path for SAHA. Results are shown in Fig. 7.
We assessed the algorithm’s performance using the Success weighted by Path Length (SPL) metric, as proposed by Anderson and Peter [42]. SPL combines binary success indication with the effectiveness of the path taken relative to the shortest possible path, given by:
@@
As shown in Table II, methods using our primitives demonstrate a higher success rate and SPL across sparse and dense environments. This implies planners with our primitives exhibit greater reliability in complex environments. In addition, compared to the pure-pursuit control law, our control law’s advantage grows with the number of obstacles increases. indicating that our control approach can improve the path tracking from primitive sets, especially in environments with dense obstacles.

### v2

@@ Abstract
Autonomous navigation across unstructured terrains, including forests and construction areas, faces unique challenges due to intricate obstacles and the element of the unknown. Lacking pre-existing maps, these scenarios necessitate a motion planning approach that combines agility with efficiency. Critically, it must also incorporate the robot’s kinematic constraints to navigate more effectively through complex environments. This work introduces a novel planning method for center-articulated vehicles (CAV), leveraging motion primitives within a receding horizon planning framework using onboard sensing. The approach commences with the offline creation of motion primitives, generated through forward simulations that reflect the distinct kinematic model of center-articulated vehicles. These primitives undergo evaluation through a heuristic-based scoring function, facilitating the selection of the most suitable path for real-time navigation. To account for disturbances, we develop a pose-stabilizing controller, tailored to the kinematic specifications of center-articulated vehicles. During experiments, our method demonstrates a $67\%$ improvement in SPL (Success Rate weighted by Path Length) performance over existing strategies. Furthermore, its efficacy was validated through real-world experiments conducted with a tree harvester vehicle - SAHA.
@@ I Introduction
Traditional planning methods [4], including sampling-based [5], search-based [6], and optimization-based strategies [7], face challenges in real-world applications, particularly within unknown environments. Those environments are characterized by frequent and unpredictable changes, making it infeasible to maintain an updated global map which sampling-based methods and grid-search methods primarily rely on. Moreover, the special and complex kinematic models of the vehicles make it difficult to find solutions efficiently through optimization-based methods in real-time operations. End-to-end learning methods, which directly map sensory observations into trajectories or actions, have gained increasing popularity due to their simplicity. However, learning-based methods struggle to incorporate strict kinematic constraints into the network, leading to potential issues in path feasibility [8].
Figure 1: Experiment of a center-articulated vehicle navigating through a construction site. A, B, and C represent three intermediate moments. The goal is set in the orange dot. The green curve shows the vehicle’s path avoiding obstacles marked by red boxes. The bottom images illustrate how obstacles influence the planning choices.
@@
This paper addresses the challenge of autonomous navigation for center-articulated vehicles by introducing an algorithm that generates motion primitives tailored to their specific kinematics, combined with a receding horizon strategy for adaptive, real-time planning and a feedback controller for path tracking within kinematic and control constraints. The main contributions of this paper are summarized as follows:
@@
    A pose-stabilizing feedback controller designed to achieve the convergence to a given target under the specific kinematic constraints and control limits of center-articulated vehicles.
@@ II Related Work
The kinematics and dynamics of CAVs have been studied [29, 30] to improve their autonomy in challenging environments, yet current planning research primarily leverages traditional methods. Nayl and Thaker’s optimization-based Model Predictive Control (MPC) approach aligns well with kinematics but lacks online efficiency [31]. Leander Peter’s use of Dubins curves achieves path traceability but lacks agility in obstacle-rich environments [32]. For these planning methods, various path-following control approaches were employed: Delrobaei and Mehdi’s Lyapunov-based approach uses polar coordinates to track the target point but does not inspect the path [33], Altafini proposes a Frenet frame-based path-tracking control [34], however, the model is complicated. Ridley and Cork’s method [35] shows the theoretical feasibility of autonomous regulation but its linearized model is constrained by the assumption of a small steer angle. A modified pure-pursuit algorithm has been proposed and successfully applied to tractors and rovers offering the advantages of simplicity and stability [36, 37], yet the steering and speed control are still coupled. While these studies address specific problems, a systematic planning and control method with efficiency and agility needed for real-time, obstacle-rich environments is yet to be seen on CAVs.
@@ III Methodology
We propose an autonomous navigation system consisting of a local path planner and a path-tracking controller. The path planner uses motion primitives to generate paths satisfying the kinematic constraints of CAVs, while the tracking controller operating at a higher frequency drives the vehicle along the planned path and accounts for disturbances.
@@ III-A Modeling of The Center-Articulated Vehicles
The schematic diagram of a CAV is shown in Fig. 2. The distances from the central joint to the centers of wheels $P_{1}$ and $P_{2}$ of the front and rear parts are $l_{1}$ and $l_{2}$ respectively. $\gamma$ is the center steering angle and the orientation angles of the front and rear parts are $\theta_{1}$ and $\theta_{2}$ respectively.
@@
We define the velocity of the center of the front part $P_{1}$ as $v$. Assuming no sideslip conditions and considering the geometric constraints at points $P_{1}$ and $P_{2}$, the relationship between the front part’s heading, velocity, steering angle, and steering velocity can be expressed as derived in [30]:
@@
$\dot{\theta}_{1}=-\frac{v_{f}\sin(\gamma)+l_{2}\dot{\gamma}}{l_{1}\cos(\gamma)+l_{2}}.$ (3)
By organizing equations (1) to (3), we derive a kinematic model with four state variables, using the front center $P_{1}$ as the virtual control point:
$\begin{bmatrix}\dot{x}_{1}\\ \dot{y}_{1}\\ \dot{\theta}_{1}\\ \dot{\gamma}\end{bmatrix}=\begin{bmatrix}\cos(\theta_{1})&0\\ \sin(\theta_{1})&0\\ -\frac{\sin(\gamma)}{L}&-\frac{l_{2}}{L}\\ 0&1\end{bmatrix}\begin{bmatrix}v\\ \dot{\gamma}\end{bmatrix},$ (4)
@@ III-B Offline Primitives Generation
The generation of primitives refers to the generation methods of state lattice primitives and control sampling primitives. According to the kinematic model in Equation (4), the vehicle’s state space includes $x_{1}$, $y_{1}$, $\theta_{1}$, $\gamma$, while the control space includes $v$ and $\dot{\gamma}$. First, we discretize the state space to create different groups of primitives. For the egocentric local planning problem, vehicle-centered coordinate frames can be used, which simplifies the vehicle’s state space as $(x_{1},y_{1})=(0,0)$ and $\theta_{1}=0$. Thus, we only discretize $\gamma$ into 30 groups, serving as the vehicle’s state lattices: $\Sigma_{i},i\in[1,30]$.
We define 15 control groups inside each state lattice $\Sigma_{i}$ as trajectories that vary from different initial control inputs. Each control group $\Sigma_{ij},(j\in[1,15])$ has different initial control inputs designed based on the mechanical limitations of vehicles and excluding significant overlap. Each control group contains trajectories that start with the same $v$ and $\dot{\gamma}$ and split twice at $t=(3/v_{j})$s and $t=(6/v_{j})$s. Trajectories in each control group are generated using forward kinematics simulation:
@@ III-C Online Receding-horizon Planner
Primitives set generated in Sec. III-B are further processed for collision detection. Drawing inspiration from Zhang’s method [9], the proposed two-step collision detection method combines offline precomputation and online path selection. Initially, potential collision points along trajectories in primitive sets are identified and mapped to grid cells, which are later filled with real-time terrain data. During online running, the planner dynamically eliminates paths based on real-time obstacle data matched to these grids, ensuring efficient and responsive navigation without excessive computational demands. As shown in Fig. 4.
@@
Figure 4: An illustration of the two-step collision detection: (a) shows an example of marking the corresponding collision grid for the Trajectory $1$. The collision body for the CAV is designed as 2 circles around the front and rear parts and an extra circle for the arm. (b) shows an online collision detection where Trajectory $3$ is marked as occluded.
@@
$S_{ij}=\frac{\sum_{k=1}^{N_{ij}}s^{ij}_{k}}{N_{ij}},$ (8)
where $N_{ij}$ is the total number of trajectory in control group $\Sigma_{ij}$. Each trajectory score $s^{ij}_{k}$ is determined by
$\begin{split}s^{ij}_{k}=&(s^{dir}+\alpha\cdot s^{dist})^{2}\cdot s^{vel}\cdot s^{state}\cdot s^{terrain}\cdot s^{p}.\end{split}$ (9)
The score terms in (9) are given by:
$s^{dist}=$ $f^{dist}\propto(D_{max}-\sqrt{dx^{2}+dy^{2}})$ (10)
$s^{dir}=$ $f^{dir}\propto{(2\pi-|d\theta}_{1}|),({2\pi-|d\theta}_{2}|)$ (11)
$s^{state}=$ $f^{state}\propto{(2\gamma_{max}-|\gamma_{k}-\gamma_{now}|)}$ (12)
$s^{vel}=$ $f^{vel}\propto{v_{k}}$ (13)
@@
Following previous methods [9], we take the coinciding part before the first split (as discussed in Sec. III-B) of the trajectories in the control group with the highest cumulative score as the optimal path. To sum up, the online local planning algorithm is described in Algorithm 1.
@@ III-D Pose-Stabilizing Control Law
The motion primitive selected by the planner provides a kinematically feasible path, but it cannot be used as a feedforward control input due to the discretization of states in the primitive generation process. Therefore, we sample a lookahead point from the selected primitive and design a variant of the pure-pursuit method inspired by Park’s method [38]. Compared to the pure-pursuit method, our controller accounts for the specialized kinematics of CAVs. It starts with the error definition in the polar coordinate. As shown in Fig. 2, the distance to the goal is $r$. Let $\phi\in(-\pi,\pi]$ represent the orientation of $P_{1}$ relative to the goal. Let $\delta\in(-\pi,\pi]$ denote the vehicle’s heading orientation relative to the lines toward the goal.
@@ III-D2 Fast Subsystem and Closed-Loop Steering
$\kappa=-\frac{1}{r}\left[k_{\delta}(\delta-\tan^{-1}(-k_{\phi}\phi))+\left(1+\frac{k_{\phi}}{1+(k_{\phi}\phi)^{2}}\right)\sin(\delta)\right]$ (22)
@@
where, $v_{op}$ is the operational speed and $v_{ax}$ is the maximum linear velocity. $\beta$ and $\lambda$ are design parameters, chosen such that $v\rightarrow v_{op}$ as $\kappa\rightarrow 0$, and $v\rightarrow v_{max}$ as $\kappa\rightarrow\kappa_{max}$.
The control law defined by Equations (23) to (25) fundamentally relies on pre-computed primitives for theoretical stability since both $\gamma$ and $\dot{\gamma}$ are mechanically constrained. When $\gamma$ or $\dot{\gamma}$ reaches the limit, Equation (23) might not be able to satisfy, resulting in potentially slow or even failure in the convergence of slow subsystem. This issue is addressed by constraining each point in the tracking paths to specifically designed primitive sets. These sets are tailored to the kinematic and mechanical constraints of center-articulated vehicles (CAVs), thereby facilitating system convergence without violating the outlined constraints.
@@ IV Experiments
We evaluate the proposed method in simulation environments and compare its performance with the baseline before carrying out real-world experiments. The test platform, SAHA (Supervised Autonomous HArvester) [39], is an autonomous tree harvester with a center-articulated chassis. The vehicle is equipped with a LiDAR for real-time environmental mapping. Key specifications of SAHA are listed in Table I. Fig. 6 depicts the actual SAHA vehicle.
@@
TABLE 3: Comparison of Motion Planning Test Results
@@ IV-A Simulation Experiments
The simulation experiment of SAHA is performed in the Gazebo simulator. We first evaluate the proposed path-following controller independently before using it in the autonomous driving pipeline. The controller is tested against the modified pure-pursuit approach, with the reference paths derived from generated primitives that accommodate different vehicle states. Specifically, three states are tested, each assessed across 90 trajectories. Cross-track error (CTE) [40] is used to evaluate the tracking error. As shown in Table 2, our method’s advantage grows with the tracking steering angle $\gamma$ increases, showing a 4.73% reduction of CTE at $\gamma=0^{\circ}$ and 27% at $\gamma=30^{\circ}$.
TABLE 2: CTE Comparison of Two Controller
@@
We then compare the proposed motion planning pipeline against the baseline method. The baseline uses a planner adapted from [9] with cropped cubic splines as primitives. These primitives, designed for searching within specified steering angles as depicted in Fig. 8, enhance the effectiveness of the search process to meet the vehicle’s kinematic constraints. The pure-pursuit (p.p.) controller is used in the baseline to track the path. In addition, we compare ours with the method that uses our primitives in the local planner, but without using our proposed pose-stabilizing controller for path following. All approaches utilize a geometric-based terrain analysis module [41] for obstacle detection without known prior maps.
@@
To evaluate robustness across various scenarios, we design four environments with varying obstacle numbers $N_{t}$ (the number of tree obstacles) and $N_{w}$ (the number of wall obstacles). For each environment, we generated 2 random maps, each spanning $40m$ by $40m$ and featuring cylinder and box obstacles to represent trees and walls. 25 target points were randomly sampled on each map, totaling 200 points across all maps. As shown in Fig. 7, the green circles are tree obstacles, gray lines are wall obstacles and the red dots are goals.
The RRT* algorithm, based on Dubins Curve, constrained by the vehicle’s steer angle and minimum turning radius, was used to find approximated shortest paths to these points with given prior maps. During our experiments, we assume after 500,000 iterations, the solution could approximate the optimal path for SAHA.
We assess the algorithm’s performance using the Success weighted by Path Length (SPL) metric, as proposed by Anderson and Peter [42]. SPL combines binary success indication with the effectiveness of the path taken relative to the shortest possible path, given by:
@@
As shown in Table 3, methods using our primitives demonstrate a higher success rate and SPL across sparse and dense environments. This implies planners with our primitives exhibit greater reliability in complex environments. In addition, compared to the pure-pursuit control law, our control law’s advantage grows with the number of obstacles increases. indicating that our control approach can improve the path tracking from primitive sets, especially in environments with dense obstacles.